# SimGRAG: Leveraging Similar Subgraphs for Knowledge Graphs Driven Retrieval-Augmented Generation

### v1

@@ A.2 Fact Verification
Appendix B Prompts
In the query-to-pattern alignment stage, Table 10 shows the prompt for MetaQA dataset, and Table 11 shows the prompt for FactKG dataset.
For verbalized subgraph-augmented generation, Table 12 shows the prompt for MetaQA dataset, and Table 13 shows the prompt for FactKG dataset.
Appendix C Implementations for Approaches
All programs are implemented with Python.
C.1 SimGRAG
C.2 Pre-trained LLMs
C.3 KG-GPT
C.4 KELP
C.5 G-Retriever
C.6 KAPING
Appendix D Detailed Error Analysis
The errors occurring during the query-to-pattern alignment step are defined as: LLM fails to follow the given instructions and examples. For example, for the query “The lady Anne Monson was born in the Darlington location of the ITL?” from FactKG dataset, the LLM gives the pattern graph with only one triple “(’Anne Monson’, ’born in’, ’Darlington’)”, which is not aligned with the query text.
The error occurred during the subgraph-augmented generation step is defined as that given the correct retrieved subgraph, the LLM fails to provide the final correct response. For example, for the question “what films did Lucky McKee star in” from the MetaQA dataset, correct subgraphs of “[(’Lucky McKee’, ’starred_actors’, ’Roman’)]” is successfully retrieved, along with the two subgraphs with lower GSD (“[(’Lucky McKee’, ’directed_by’, ’All Cheerleaders Die’)]” and “[(’Lucky McKee’, ’directed_by’, ’May’)]”). However, the LLM gives the final response of “According to the evidences, there is no direct connection between Lucky McKee and a film they starred in. The graphs only mention that Lucky McKee directed films (’All Cheerleaders Die’ and ’May’), but do not provide information about the films they acted in.”
Errors occurring during the pattern-to-subgraph alignment phase are defined as: LLM follows the given instructions and examples to generate a satisfactory pattern graph, but the retrieval algorithm fails to retrieve the ground-truth subgraph for the query. It is because the ground-truth subgraphs have different structures and thus cannot be successfully aligned with the ground-truth subgraphs. For example, for the query “A food is classed as a Dessert and can be served warm (freshly baked) or cold.”, the LLM-generated pattern graph is “[(’UNKNOWN food 1’, ’classed as’, ’Dessert’), (’UNKNOWN food 1’, ’served’, ’"warm"’), (’UNKNOWN food 1’, ’served’, ’"cold"’)]”. However, the ground-truth subgraphs have the structure like “[(’The food name’, ’classed as’, ’Dessert’), (’The food name’, ’served’, ’"warm (freshly baked) or cold"’)]”.
Appendix E Parameters for Grid Search
We conduct the grid search for evaluating the top-$k$ retrieval algorithm and its optimized one on the FactKG dataset using DBpedia knowledge graph. We fix $k=1$ and try all combinations of the other parameters $k^{(n)}\in\{128,256,512,1024,2048,4096,8192,16384\}$, $k^{(r)}\in\{128,256,512\}$, $k^{(t)}\in\{1,2,4,8,16\}$. For 100 queries, any program run out of the time limit of 10,000 seconds will be terminated and not reported. In Figure 5, the point at retrieval Hits@1=1.0 is achieved by using $k^{(n)}=16384$, $k^{(r)}=512$ and $k^{(t)}=16$.
@@ Appendix E Parameters for Grid Search
Table 10: The query-to-pattern alignment prompt used in MetaQA dataset.
@@
Table 12: The verbalized subgraph-augemented generation prompt used in MetaQA dataset.

### v2

@@ A.2 Fact Verification
Head Relation Tail
Berlin country Germany
United States governmentType Republic
Harry Potter author J. K. Rowling
Albert Einstein award Nobel Prize in Physics
Terrance Shaw college Stephen F. Austin State University
Association for the Advancement of Artificial Intelligence type Scientific society
Nvidia industry Computer hardware
Table 14: Examples triples in the DBpedia knowledge graph used for FactKG dataset.
                             Example statements
                             1. It was Romano Prodi who was the prime minister.
                             2. Are you familiar with Terrance Shaw? He also attended college.
                             3. Yes, Anastasio J. Ortiz was the Vice President.
Table 15: Example statements from the FactKG dataset.
@@
Table 16: The query-to-pattern alignment prompt used for KGQA task.
Appendix B Prompts
For query-to-pattern alignment, Table 16 shows the prompt for KGQA tasks, including MetaQA, PathQuestions and WC2014 datasets. Table 17 shows the prompt for the fact verification task, i.e., FactKG dataset.
For verbalized subgraph-augmented generation, Table 18 shows the prompt for KGQA tasks, including MetaQA, PathQuestions and WC2014 datasets. Table 19 shows the prompt for the fact verification task, i.e., FactKG dataset.
At each step for processing each dataset, our prompts utilize exactly the same guidances and few-shot examples across different query pattern structures. For example, all 1/2/3-hop queries in the MetaQA dataset share the identical prompt at the step of query-to-pattern alignment.
@@ Appendix B Prompts
Table 18: The verbalized subgraph-augemented generation prompt used for KGQA task.
@@
Appendix C Implementations for Approaches
All programs are implemented with Python.
C.1 SimGRAG
C.2 Pre-trained LLMs
C.3 KG-GPT
C.4 KELP
C.5 G-Retriever
C.6 KAPING
Appendix D Discussion about Oracle Entities
As discussed in Section 1, in real applications, users might not always know the precise entity IDs related to their query. Thus, an ideal approach should not require users to specify the oracle entities. However, both KG-GPT Kim et al. (2023a) and KELP Liu et al. (2024) expand subgraphs or paths from the user-provided oracle entities, while G-Retriever He et al. (2024) restricts the KG to a 2-hop oracle entity neighborhood. In other words, they need to know which entities are exactly correct before running, and the search space will be constrained in the ground truth area, thereby reducing the problem hardness.
Though all methods will work better with the oracle entities, experimental evaluation in Table 1 shows that even when we allow certain baselines to benefit from using the oracle entities, their performance still underperforms the SimGRAG method that does not require such entities. In other words, if we do not provide them for such baselines, their performance may degrade further.
Moreover, it is intuitive to use the results of a top-$k$ entity linker as a substitute for the oracle entity in certain baselines. However, it would significantly increase the computational complexity and latency, since these methods might need to run the entire pipeline independently for each candidate entity. In contrast, the SimGRAG method naturally avoids relying oracle entities without such independent redundant computations.
Furthermore, unlike existing approaches, the internal mechanism of SimGRAG method is designed to better handle and filter out those noisy entities in real-world KGs. As discussed in Section 4.2, we use the Graph Semantic Distance (GSD) metric, which can effectively incorporate somewhat distant entities or relations that still contribute valuable evidence to the overall subgraph. For example, Figure 4 shows that a candidate entity whose semantic distance is ranked $112$ can still be used in the retrieved subgraph.
Generally, the proposed SimGRAG method is closer to the ideal feature through a carefully designed mechanism, while ensuring better performance than the baseline methods.
Appendix E Detailed Error Analysis
The errors occurring during the query-to-pattern alignment step are defined as: LLM fails to follow the given instructions and examples. For example, for the query “The lady Anne Monson was born in the Darlington location of the ITL?” from FactKG dataset, the LLM gives the pattern graph with only one triple “(‘Anne Monson’, ‘born in’, ‘Darlington’)”, which is not aligned with the query text.
The error occurred during the subgraph-augmented generation step is defined as that given the correct retrieved subgraph, the LLM fails to provide the final correct response. For example, for the question “what films did Lucky McKee star in” from the MetaQA dataset, correct subgraphs of “[(‘Lucky McKee’, ‘starred_actors’, ‘Roman’)]” is successfully retrieved, along with the two subgraphs with lower GSD (“[(‘Lucky McKee’, ‘directed_by’, ‘All Cheerleaders Die’)]” and “[(‘Lucky McKee’, ‘directed_by’, ‘May’)]”). However, the LLM gives the final response of “According to the evidences, there is no direct connection between Lucky McKee and a film they starred in. The graphs only mention that Lucky McKee directed films (‘All Cheerleaders Die’ and ‘May’), but do not provide information about the films they acted in.”
Errors occurring during the pattern-to-subgraph alignment phase are defined as: LLM follows the given instructions and examples to generate a satisfactory pattern graph, but the retrieval algorithm fails to retrieve the ground-truth subgraph for the query. It is because the ground-truth subgraphs have different structures and thus cannot be successfully aligned with the ground-truth subgraphs. For example, for the query “A food is classed as a Dessert and can be served warm (freshly baked) or cold.”, the LLM-generated pattern graph is “[(‘UNKNOWN food 1’, ‘classed as’, ‘Dessert’), (‘UNKNOWN food 1’, ‘served’, ‘"warm"’), (‘UNKNOWN food 1’, ‘served’, ‘"cold"’)]”. However, the ground-truth subgraphs have the structure like “[(‘The food name’, ‘classed as’, ‘Dessert’), (‘The food name’, ‘served’, ‘"warm (freshly baked) or cold"’)]”.
Appendix F Query Pattern Structures
Following the previous study Zhang et al. (2016), all queries (i.e., the pattern structure) in our experiments can be categorized into the following six types. For simplicity, we focus on topological structures and ignore the edge directions.
  • 1-hop Path: Find an edge from a known subject $s$ to another entity $e$.
  • 2-hop Path: Find a 2-hop path from a known subject $s$ to another known or unknown entity $e$.
  • 3-hop Path: Find a 3-hop path from a known subject $s$ to another known or unknown entity $e$.
  • 2-hop Conjunction: Find two distinct edges that link two known subjects $s_{1}$ and $s_{2}$ with an unknown entity $e$.
  • 3-hop Conjunction: Find an edge that links an known subject $s_{1}$ with an unknown entity $e$, as well as a 2-hop path that connects another known subjects $s_{2}$ with the same entity $e$.
  • 3-hop Star: Find three distinct edges that links to the same known or unknown entities $e$.
Appendix G Parameters for Grid Search
We conduct the grid search for evaluating the top-$k$ retrieval algorithm and its optimized one on the FactKG dataset using DBpedia knowledge graph. Specifically, we randomly sample 100 queries that correctly generate patterns and manually identify the ground truth subgraphs for each query to evaluate retrieval performance using retrieval Hits@1. We fix $k=1$ and try all combinations of the other parameters $k^{(n)}\in\{128,256,512,1024,2048,4096,8192,16384\}$, $k^{(r)}\in\{128,256,512\}$, $k^{(t)}\in\{1,2,4,8,16\}$. For 100 queries, any program run out of the time limit of 10,000 seconds will be terminated and not reported. In Figure 5, the point at retrieval Hits@1=1.0 is achieved by using $k^{(n)}=16384$, $k^{(r)}=512$ and $k^{(t)}=16$.
Appendix H Experiments on WebQSP Dataset
We also test SimGRAG on the WebQuestionSP (WebQSP) dataset Yih et al. (2016) using WikiData Vrandečić and Krötzsch (2014), which is the most popular and active KG. We use the 2015 WikiData dump to align with WebQSP’s creation time, which contains 26 million nodes and 57 million edges. We use the WebQSP-WD test set Sorokin and Gurevych (2018), a corrected version of the original WebQSP dataset for WikiData compatibility.
Since Sorokin and Gurevych (2018) mentioned that not all queries are guaranteed answerable with WikiData, we manually verified each query in the test set, excluding those without any supporting evidence in WikiData. Specifically, each question in the WebQSP-WD test set is associated with a set of topic entities and answer entities, which were mapped from their original Freebase IDs to WikiData IDs. Since WebQSP dataset is all about questions within 2-hops He et al. (2024), for each question, we extracted the 2-hop neighborhood of the topic entities, as well as the 2-hop neighborhood of the answer entities in WikiData KG, and union all these edges together to form a single subgraph. To determine whether a question is supported by WikiData, we compared the question with its corresponding merged subgraph. A question was considered unsupported if either of the following conditions held:
  • None of the topic entities are connected to any answer entities within the subgraph.
  • All connections between the topic entities and the answer entities are completely irrelevant to the intent of the question. Note that we never require the subgraph to contain a path that directly answers the question. As long as a human could infer the correct answer by reasoning over the whole connected subgraph, we considered the question to be supported.
After manually ensuring the quality and reliability of the test set, the proposed SimGRAG method achieves the Hits@1 of 87.7%.